The following two reactions are important in the blast furnace production of iron metal from iron ore (Fe2O3): 2C(s) + O2(g) → 2CO(g) Fe2O3(s) + 3CO(g) → 2Fe(s) + 3CO2(g) Using these balanced reaction, how many moles of O2 are required for the production of 5.35 kg of Fe?


1
Expert's answer
2021-06-21T05:31:31-0400

1 mole of O2 is needed to produce 2 moles of CO

3moles of CO, 1.5moles of O2 is needed

for 2moles of Fe, 1.5 moles of O2 is needed

for 1mole of Fe, 0.75 moles of O2 is required

moles of Fe=5350/55.8=95.88 moles

therefore the moles of O2 needed =95.88*0.75 =71.97 moles